CaCl2(aq) + Na2CO3(aq) --> CaCO3(s) + 2NaCl(aq)

20 mL of 0.5M CaCl2 is added to 10 mL of 1.5M Na2CO3. Determine the number of moles of CaCl2 and Na2CO3 and calculate the the number of moles of CaCO3 that would be expected. In and experiment you produced 0.898g of CaCO3 what is the precent yeild?

% yeild=________________
Answer #1

CaCl2(aq) + Na2CO3(aq) --> CaCO3(s) + 2NaCl(aq)

20 mL of 0.5M CaCl2 is added to 10 mL of 1.5M Na2CO3.

1) Determine the number of moles of CaCl2 and Na2CO3

Molarity = no. of moles/Volume of solution(L)

a) 20 mL= 0.02 L of 0.5M CaCl2 ; no. of moles = M*V(L) = 0.5*0.02 = 0.01 moles of CaCl2

b) 10 mL=0.01 L of 1.5M Na2CO3 ; no. of moles = M*V(L) = 1.5*0.01 = 0.015 moles of Na2CO3

2) calculate the the number of moles of CaCO3 that would be expected.

From balanced equation, 1mole of CaCl2 gives 1 moles of CaCO3. Thus 0.01 mole of CaCl2 gives 0.01 mole of CaCO3

amount of CaCO3 will be formed = moles*molar mass=0.01*101 = 1.01 gms (theoritical yield)

3) In and experiment you produced 0.898g of CaCO3 what is the precent yeild?

%yield = (actual yield/ theoritical yield)*100

%yield = (0.898/1.01)*100 =88.91 %

% yeild= 88.91%